OpenStudy (anonymous):

the height of a triangle is 4 inches less than its base. Find the base and height of the triangle if its area is 70 square inches.

OpenStudy (anonymous):



OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

height = h , base = h + 4 area = (1/2) * h (h + 4) = 70 1/2 * h^2 + 4h = 70 h^2 + 4h = 140 h^2 + 4h - 140 = 0 h = 10 and base = 14

OpenStudy (anonymous):

Area of triangle = 1/2 base * height. Base = b => area = 1/2 * (b-4) * b = 1/2 b^2 -2b =70. Thus b^2 - 4b = 140. Solve quadratic equation or spot b = 14.